OpenStudy (anonymous):

The quadratic function f(x) = x^2-Kx+2K has a vertex whose y-coordinate is 5. Find all the possible values of K.

OpenStudy (tkhunny):

f(x) = (x^2 - Kx + _____) + 2k - _____ We're just Completing the Square. Fill in the two blanks with the SAME correct expression.

OpenStudy (anonymous):

Completing the square is where you take the middle term and divide it by 2.

OpenStudy (abmon98):

b^2-4ac (k)^2-4(1)(2k-5) k^2-8k+40 solve for k

OpenStudy (anonymous):

so you did it with the quadratic formula?

OpenStudy (tkhunny):

Why are we doing all that? Just complete the square!

OpenStudy (anonymous):

-1 +- sqrt(1^2-4(1)(2)/2(1)

OpenStudy (anonymous):

I don't know how to complete the square...

OpenStudy (anonymous):

I have used the quadratic formula almost every time

OpenStudy (tkhunny):

f(x) = (x^2 - Kx + (K/2)^2) + 2K - (K/2)^2 The y-coordinate of the Vertex is 2K - (K/2)^2 Solve: 2K - (K/2)^2 = 5

OpenStudy (anonymous):

K - (K/2)^2 = 5/2

OpenStudy (anonymous):

K - K = 5/2

OpenStudy (anonymous):

So I divide 2 by all the terms and then square root both sides

OpenStudy (tkhunny):

It's a quadratic equation. You've solved dozens of these. 2K - (K/2)^2 = 5 2K - (K^2)/4 = 5 8K - K^2 = 20 K^2 - 8K + 20 = 0 Abmon98 had the right idea, but also a little error there, right at the end.

OpenStudy (anonymous):

So now I need to solve for the K values?

OpenStudy (anonymous):

It is in quadratic form

OpenStudy (anonymous):

It doesn't factor and has no real solutions